Two charges Q1=+3.5 nC and Q2=-5.3 nC are separated by 10 cm. Calculate the electric potential at point B

1
Expert's answer
2022-03-02T14:42:15-0500

Electric potential

"V_1=\frac{kq_1}{r}"

"V_1=\frac{9\times10^9\times3.5\times10^{-9}}{5\times10^{-2}}=630V"

"V_2=\frac{kq_2}{r}"

"V_2=-\frac{9\times10^9\times5.3\times10^{-9}}{5\times(10)^{-2}}=-954V"

"V_{net}=V_1+V_2"

"V_{net}=630-954=-324V"